# “It’s not a representation of me”: Examining Accent Bias and Digital Exclusion in Synthetic AI Voice Services

### v1

@@ 8.1. Ethical Considerations
Our study was reviewed and approved by our university’s IRB (protocol # redacted for blind reviewing). This included strategies for mitigating adverse impacts, procedures for incident reporting when necessary, and ensuring all authors completed and maintained an up-to-date CITI Program training certificate. All crowdworkers were compensated at a rate of \$15 USD per hour, while interview participants received \$20 USD gift cards as compensation for their participation. Participation was voluntary and there were no consequences if participants chose to withdraw, but no participant did so. All participants provided informed consent and we encouraged questions throughout the interview study for clarity and comfort. The collected data were only accessible to the authors.
@@ 8.2. Adverse and Unintended Impact
Our primary adverse impact concern is unintentionally advertising these voice clone tool’s potential to fuel impersonation scams. This is highlighted by P25’s statement on the utility of the voice clone tools they were presented in the interview, stating “…I can definitely think of use cases, but like none that I’d participate in myself like, unless I became like a con artist…” Secondly, despite having technical justifications to do so, we may inadvertently contribute to perpetuating overgeneralized and reductive taxonomies of regional accents, including the continent-wide category of “African” and subcontinent-wide category of “Indian”. Such classifications can undermine regional diversity and may unintentionally validate problematic or colonial-era linguistic hierarchies. Finally, the study’s findings, while specific to Speechify and ElevenLabs, might be extrapolated by others to represent all AI voice technologies, leading to an inaccurate understanding of the broader landscape. We strongly encourage future researchers to adapt their studies to the nuances of the specific systems that they choose to study.
@@ 8.3. Author Positionality
[This section is redacted for the double-blind submission.]
@@ A.1. Speechify and ElevenLabs
(a) Speechify
(b) ElevenLabs
@@
The left screenshot is the user interface of Speechify. This screenshot includes options for the user to select what language to generated text in and options for stock synthetic AI voices. The right screenshot is ElevenLab’s user interface. This interface also has a drop down menu of what stock synthetic AI voice users can choose from as well as the side menu showcasing the services’ other technical capabilities such as voice changer, voiceover studio and dubbing studio. .
@@ A.2. Prolific Crowdworkers’ Demographics
(a) Prolific Crowdworkers’ Self-reported Gender (Study 1)
(b) Prolific Crowdworkers’ Self-reported Age (Study 1)
(c) Prolific Crowdworkers’ Self-reported Gender (Study 2)
(d) Prolific Crowdworkers’ Self-reported Age (Study 2)
Figure 11. Self-reported gender and age from the studies detailed in § 4.1 and § 4.2.
@@
The top-left subfigure is a bar chart depicting the proportions of Prolific crowdworkers by their self-reported gender for Study 1. The distribution shows the majority identify as male or female. The top-right subfigure is a bar chart illustrating the proportions of Prolific crowdworkers by their self-reported age for Study 1. Most participants are under the age of 60. The bottom-left subfigure is a bar chart depicting the proportions of Prolific crowdworkers by their self-reported gender for Study 2. The distribution shows the majority identify as male or female. The bottom-right subfigure is a bar chart illustrating the proportions of Prolific crowdworkers by their self-reported age for Study 2. Most participants are under the age of 60.
@@ A.5. Mean Opinion Score (MOS) When Accent is Predicted Correctly
Table 2. The Mean Opinion Score (MOS) regression model results when accent was predicted correctly. Reference levels are in comparison to the American Accent, ElevenLabs, and high-pitched voices.
A.6. Voice Quality Analysis Across Pitch and Service
@@
The left subfigure presents a Sankey plot that visualizes pitch alignment between Prolific crowdworkers and Speechify’s pitched voice labels. The percentages of agreed classifications across the two accents are: Low (99.28%) and High (99.64%). The right subfigure presents a similar Sanky plot but with pitch labels from ElevenLabs: Low (83.66%) and High (98.35%)
@@
A bar chart comparing the Mean Opinion Score (MOS) for Speechify and ElevenLabs. The x-axis represents the MOS, ranging from 3 to 4.5, while the y-axis lists the two services. Two bars represent the the two pitched voices, each with 95% confidence intervals. A dotted red line at 4.4 marks the baseline for excellent quality, but none of the bars reach or exceed this value.
A.7. Interview Questions
The following are the semi-structured interview questions for both in-person and remote sessions. Participants were assigned a sequential ID at the time the interview was conducted.
A.7.1. Demographic and Accent Mimicry Quality (AMQ) Questionnaires
Figure 17. Sample survey participants completed, providing demographic information and evaluations of each cloned version of their voice. These evaluations followed the Accent Mimicry Questionnaire (AMQ) measure and detailed in § 4.3.
A.7.2. Voice and Accent Perception Questions
@@
A.7.3. AI Speech Technology Familiarity
Are you familiar with AI speech generating services? Have you heard of such services? Have you used any of these services before? If yes, how so? In what contexts do you think AI speech generating services are being used for? Would you use it for any of those purposes?
A.7.4. Listening Experience Follow-up Questions (after each pair)
How did you feel while listening to the generated voice clips (e.g. feelings of surprise, satisfaction, discomfort, etc.)? Do you feel like they represented you? How accurate do you think the generated voice clips captured your accent and speech patterns? (overall and for each service) In what way have the voice clips accurately reflected your identity and how you communicate? Would you be comfortable using a voice cloning service like this one to read a script for a podcast or other long-form spoken content of your voice? What concerns do you have about using these types of services? Do you have any authenticity, privacy, or ethical considerations? What concerns do you have for others about using these types of services?
A.8. Intimate Relationship with Voice and Accent
@@ A.8.1. Experiences of Othering
A.8.2. Voice and Accent are Part of Identity
@@
A.8.3. Voice and Accent are Culture and Class Identifiers
@@
A.8.4. Code-switching as a Strategy

### v2

@@ 8.1. Ethical Considerations
Our study was reviewed and approved by Northeastern University’s IRB (Protocol #2024-05-13). This included strategies for mitigating adverse impacts, procedures for incident reporting when necessary, and ensuring all authors completed and maintained an up-to-date CITI Program training certificate. All crowdworkers were compensated at a rate of \$15 USD per hour, while interview participants received \$20 USD gift cards as compensation for their participation. Participation was voluntary and there were no consequences if participants chose to withdraw, but no participant did so. All participants provided informed consent and we encouraged questions throughout the interview study for clarity and comfort. The collected data were only accessible to the authors.
@@ 8.2. Adverse and Unintended Impact
Our primary adverse impact concern is unintentionally advertising these voice clone tool’s potential to fuel impersonation scams. This is highlighted by P25’s statement on the utility of the voice clone tools they were presented in the interview, stating “…I can definitely think of use cases, but like none that I’d participate in myself like, unless I became like a con artist…” Secondly, despite having technical justifications to do so, we may inadvertently contribute to perpetuating overgeneralized and reductive taxonomies of regional accents, including the continent-wide category of “African” and subcontinent-wide category of “Indian”. Such classifications can undermine regional diversity and may unintentionally validate problematic or colonial-era linguistic hierarchies. To counter this, we encourage following the #BenderRule when naming the language or accent of study (Bender, 2019). Third, by focusing exclusively on English, our study risks perpetuating the overrepresentation of English in language technologies, potentially marginalizing speakers of less-resourced languages. Finally, the study’s findings, while specific to Speechify and ElevenLabs, might be extrapolated by others to represent all AI voice technologies, leading to an inaccurate understanding of the broader landscape. We strongly encourage future researchers to adapt their studies to the nuances of the specific systems that they choose to study.
@@ 8.3. Author Positionality
At the time of conducting this research, all authors were affiliated with academic institutions or research organizations focused on AI and machine learning, fairness, human-computer interaction, privacy, and/or security. This study is rooted in our interest in understanding how these technologies intersect with linguistic diversity and lived experiences of users. We recognize how transformative the potential of synthetic AI speech generation and voice cloning technologies are as well as their broader societal implications. Our experiences and perspectives informs our research design, analysis, and interpretation, with the goal of advocating for advancing equitable and inclusive AI speech technologies. The first, second, and third authors have high-pitched American accents. The fourth and fifth authors have low-pitched American accents while the last author has a low-pitched Indian Bengali accent.
Acknowledgements.
This research was supported in part by a Northwestern CASMI grant. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the funders. We would also like to thank all the participants who shared their time, insights, and experiences during the interviews, and to the anonymous reviewers for their feedback during the peer review process.
@@ A.1. Speechify and ElevenLabs
Figure 10. Speechify’s user interfaces.
[…]
The screenshot is the user interface of Speechify. This screenshot includes options for the user to select what language to generated text in and options for stock synthetic AI voices.
@@
The screenshot is ElevenLab’s user interface. This interface also has a drop down menu of what stock synthetic AI voice users can choose from as well as the side menu showcasing the services’ other technical capabilities such as voice changer, voiceover studio and dubbing studio.
@@ A.2. Prolific Crowdworkers’ Demographics
Figure 12. Prolific Crowdworkers’ self-reported gender (Study 1)
@@
The bar chart depicts the proportions of Prolific crowdworkers by their self-reported gender for Study 1. The distribution shows the majority identify as male or female.
Figure 13. Prolific Crowdworkers’ self-reported age (Study 1)
[…]
The bar chart illustrates the proportions of Prolific crowdworkers by their self-reported age for Study 1. Most participants are under the age of 60.
Figure 14. Prolific Crowdworkers’ self-reported gender (Study 2)
[…]
The bar chart depicts the proportions of Prolific crowdworkers by their self-reported gender for Study 2. The distribution shows the majority identify as male or female.
Figure 15. Prolific Crowdworkers’ self-reported age (Study 2)
[…]
The bar chart illustrates the proportions of Prolific crowdworkers by their self-reported age for Study 2. Most participants are under the age of 60.
@@ A.5. Mean Opinion Score (MOS) When Accent is Predicted Correctly
Figure 19. The Mean Opinion Score (MOS) regression model results when accent was predicted correctly. Reference levels are in comparison to the American Accent, ElevenLabs, and high-pitched voices.
A.6. Mean Opinion Score (MOS) Regressions by Service
Estimate (Std. Error)
(Intercept) $4.09$ $(0.20)^{***}$
African Accent $0.02$ $(0.26)$
Australian Accent $-0.04$ $(0.26)$
British Accent $-0.004$ $(0.26)$
Indian Accent $0.15$ $(0.29)$
Low-pitched Voice $-0.39$ $(0.16)^{*}$
${}^{***}p<0.001$; ${}^{**}p<0.01$; ${}^{*}p<0.05$
Figure 20. The Mean Opinion Score (MOS) regression model results from Speechify. Reference levels are in comparison to the American Accent and high-pitched voices.
Estimate (Std. Error)
(Intercept) $4.09$ $(0.10)^{***}$
African Accent $0.28$ $(0.14)^{*}$
Australian Accent $0.31$ $(0.14)^{*}$
British Accent $0.33$ $(0.14)^{*}$
Indian Accent $0.07$ $(0.14)$
Low-pitched Voice $-0.21$ $(0.08)^{*}$
${}^{***}p<0.001$; ${}^{**}p<0.01$; ${}^{*}p<0.05$
Figure 21. The Mean Opinion Score (MOS) regression model results from ElevenLabs. Reference levels are in comparison to the American Accent and high-pitched voices.
A.7. Voice Quality Analysis Across Pitch and Service
@@
The left subfigure presents a Sankey plot that visualizes pitch alignment between Prolific crowdworkers and Speechify’s pitched voice labels. The percentages of agreed classifications across the two accents are: Low (99.28%) and High (99.64%). The right subfigure presents a similar Sanky plot but with pitch labels from ElevenLabs: Low (83.66%) and High (98.35%).
@@
A bar chart comparing the Mean Opinion Score (MOS) for Speechify and ElevenLabs. The x-axis represents the MOS, ranging from 3 to 4.5, while the y-axis lists the two services. Two bars represent the two pitched voices, each with 95% confidence intervals. A dotted red line at 4.4 marks the baseline for excellent quality, but none of the bars reach or exceed this value.
A.8. Interview Questions
The following are the semi-structured interview questions for both in-person and remote sessions. See Figure 24 for the sample survey questions and Figure 25 for the semi-structured interview questions. Participants were assigned a sequential ID at the time the interview was conducted.
A.8.1. Demographic and Accent Mimicry Quality (AMQ) Questionnaires
Figure 24. Sample survey participants completed, providing demographic information and evaluations of each cloned version of their voice. These evaluations followed the Accent Mimicry Questionnaire (AMQ) measure and detailed in § 4.3
A.8.2. Voice and Accent Perception Questions
@@
A.8.3. AI Speech Technology Familiarity
Are you familiar with AI speech generating services? Have you heard of such services? Have you used any of these services before? If yes, how so? In what contexts do you think AI speech generating services are being used? Would you use it for any of those purposes?
A.8.4. Listening Experience Follow-up Questions (after each pair)
How did you feel while listening to the generated voice clips (e.g. surprise, satisfaction, discomfort)? Do you feel like they represented you? How accurate do you think the generated clips captured your accent and speech patterns (overall and per service)? In what way have the clips reflected your identity and communication style? Would you use a voice cloning service like this for podcasts or long-form audio content? What concerns do you have about using these services—authenticity, privacy, or ethical issues? What concerns do you have for others using these services?
Figure 25. Semi-structured interview questions covering voice perception, AI familiarity, and reactions to generated voice samples.
A.9. Intimate Relationship with Voice and Accent
@@ A.9.1. Experiences of Othering
A.9.2. Voice and Accent are Part of Identity
@@
A.9.3. Voice and Accent are Culture and Class Identifiers
@@
A.9.4. Code-switching as a Strategy